OpenStudy (anonymous):

A net force of magnitude 36N gives an object of mass m1 an acceleration of 6.0 m/s/s. The same net force gives m1 and another object of mass m2 fastened together an acceleration of 2.0 m/s/s. What acceleration will m2 experience if the same net force acts on it alone?

OpenStudy (mos1635):

aply newtons 2nd low to m1 m1=.... aply newtons 2nd low to m1 and m2 (together (m1+m2) is that egnaph to get you started?

OpenStudy (anonymous):

first u have to find m1 u are given F=36N, a=6m/s^2 F=ma m1=6kg now in second part u are said if both are together and net force is same F=a(m1 + m2) m2=12kg now in third part u are said if force is same find acceleration F=m2a a=3m/s^2